OpenStudy (anonymous):

Please help! Calculus questions 1.) Find the critical numbers of f(x)=(x-3)^1/3 2.) Can someone check my work I keep getting a really large number and it makes me think i messed up somewhere. The question is find the points of inflection and identify where the graph of f(x)=-x^4+24x^3 is concave up or concave down. f(x)=-x^4+24x^3 f'(x)=-4x^3+72x^2 f"(x)=-12x^2+144x=0 -12x^+144x=0 12x(-x+12)=0 x=0 and x=12 are the critical #'s so i tested -1,1,11,13 f" f"(-1)=-156, f"(1)=132, f"(11)=132, f"(13)=-156 then used the original f(x) plugged in 0 and 12 and got (0,0) and (12,20736)

OpenStudy (anonymous):

\[f(x)=(x-3)^{1/3}~~\Rightarrow~~f'(x)=\frac{1}{3}(x-3)^{-2/3}\] Critical points occur for \(x\) that make \(f'(x)\) zero or undefined, in this case \(x=3\).

OpenStudy (anonymous):

For the second question: inflection points are like critical points for a function's derivative. \[\begin{align*}f(x)&=-x^4+24x^3\\ f'(x)&=-4x^3+72x^2\\ f''(x)&=-12x^2+144x \end{align*}\] Possible inflection points occurs when \(f''(x)\) is zero or undefined. Since \(f''(x)\) is a polynomial, you don't have to worry about the undefined case. \[\begin{align*}-12x^2+144x&=0\\ x^2-12x&=0\\ x(x-12)&=0\\ x&=0,12 \end{align*}\] Now you check the intervals: \[\begin{matrix} \text{interval}&&\text{test point}&&\text{sign of }f''(\text{test point})\\ (-\infty,0)&&-1&&-\\ (0,12)&&1&&+\\ (12,\infty)&&13&&- \end{matrix}~~\iff~~\begin{matrix}.\\\text{concave down}\\\text{concave up}\\\text{concave down}\end{matrix}\] Notice that you only need one test point from each interval - the sign of the derivative will be the same to the right of 0 as it is to the left of 12. So you have actual inflection points at \(x=0\) and \(x=12\), which give \((0,0)\) and \((12,20736)\). Your work is correct. Don't let big numbers make you doubt yourself. It makes sense that you get such a big number, given that you have a quartic polynomial. You should expect big numbers the further you get from the original polynomial's roots (which were \(x=0\) and \(x=24\)).

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

You're welcome!